OpenStudy (anonymous):

5(5x + -3) = 5(3x + 7)

OpenStudy (anonymous):

Distribute 5 on both sides 25x-15= 15x+35 combine like terms, i would subtract 15 so you wont get a negative- 10x-15=35 Add 15 to both sides- 10x=50 Divide each side by 10 to get x alone- x=5 & if you go back and plug x=10 in, they both come out to 110.

OpenStudy (anonymous):

Plug x=5 not x=10, sorry

OpenStudy (anonymous):

y would you do that!!! lol i liked it better the other way i understand it better

OpenStudy (anonymous):

5(5x-3)=5(3x+7) 25x-15=15x+35 10x-15=35 10x=50 x=5

OpenStudy (anonymous):

3(-9n + 5) + -4 = -16 -27n+15 +-4=-16 -27n+11=-16 -27n=-27 n=1 can you check if this is rite

OpenStudy (anonymous):

@pinkberry501

OpenStudy (anonymous):

Sure, -27n+15-4=-16 3(-9(1)+5)-4=-16 -27n+11=-16 3(-9+5)-4=-16 -27n=-27 3(-4)-4=-16 n=1 -12-4=-16 -16=-16 Yes that is right!

OpenStudy (anonymous):

@pinkberry501 thanks!! :) :) :) :)